OpenStudy (anonymous):

Estimate 3 and 5/7 times 4 and 3/4 by rounding each factor to the nearest whole number. Explain your choices for rounding.

OpenStudy (anonymous):

3 5/7 should be what when you round it to the nearest whole number? Same with 4 3/4.

OpenStudy (hoblos):

3 5/7 = 26/7= 3.714... so we round it to 4 4 3/4 =19/4= 4.75 so we round it to 5 =>4*5=20 i guess so

OpenStudy (anonymous):

You really don't need to convert into decimals. 5/7 is clearly higher than 1/2. So, it rounds to next higher whole number. Same with 3/4. Anyway, writing out answers to people is a terrible idea on this forum. Explaining how to get to the answer is more helpful.

OpenStudy (hoblos):

people here have different minds, some need explanations and others need formulas and ways and there is some who want only the final answer so we don't know what to do